OpenStudy (precal):

How do I convert .096 radian per seconds to 5.5 degrees per second

OpenStudy (anonymous):

\[\frac{.096\text{ rad}}{1\text{ s}}\times\frac{180\text{ deg}}{\pi\text{ rad}}=\cdots\]

OpenStudy (precal):

thank you

OpenStudy (anonymous):

yw